OpenStudy (anonymous):

NEED THIS DONE BY MIDNIGHT PLEASE HELP!? Each child ticket for a ride costs $3, while each adult ticket costs $5. If the ride collected a total of $115, and 33 tickets were sold, how many of each type of ticket were sold?

OpenStudy (anonymous):

Is it multiple choice?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Plug in each choice until one matches up

OpenStudy (anonymous):

That's how I do those problems when in doubt :P

OpenStudy (anonymous):

ok thanks i'll try

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

8 child and 25 adult 17 child and 16 adult 156 child and 41 adult 25 child and 8 adult

OpenStudy (anonymous):

3c+5a=115

OpenStudy (jdoe0001):

\(\large \begin{array}{lcccllll} &\textit{tickets sold}&\textit{price for each}&\textit{total for all}\\ \hline\\ children&c&3&3c\\ adults&a&5&5c\\ \hline\\ &c+a=33&&3c+5c=115 \end{array}\) if you were to solve say c+a=33 for "a", what would you get?

OpenStudy (jdoe0001):

hmmm I even have a typo... hmm ok well \(\large \begin{array}{llll} &\textit{tickets sold}&\textit{price for each}&\textit{total for all}\\ \hline\\ children&c&3&3c\\ adults&a&5&5c\\ \hline\\ &c+a=33&&3c+5a=115 \end{array}\)

OpenStudy (jdoe0001):

yet anotehr typo shoot \(\large \begin{array}{lccclll} &\textit{tickets sold}&\textit{price for each}&\textit{total for all}\\ \hline\\ children&c&3&3c\\ adults&a&5&5a\\ \hline\\ &c+a=33&&3c+5a=115 \end{array}\) anyhow, solve the 1st equation c+a=33 by "a" first

OpenStudy (anonymous):

ohh ok i understand now!

OpenStudy (anonymous):

thank you so much!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

do you have time for more?

OpenStudy (jdoe0001):

easier to post anew... gives more exposure, thus more answers and we can also revise each other :)